OpenStudy (venomblast):

cos(9x)=1/2 how to get cosx by itself

OpenStudy (anonymous):

cos(a+b)=cos(a)cos(b)-sin(a)sin(b) sin(a+b)=sin(a)cos(b)+cos(a)sin(b) cos(x+8x)=cos(x)cos(8x)-sin(x)sin(8x) cos(4x+4x)=cos(4x)cos(4x)-sin(4x)sin(4x) sin(4x+4x)=2sin(4x)cos(4x) continue simplifying until cos(9x) is left in terms of cos(x), sin(x) then use the identity sin(x)^2+cos(x)^2=1

OpenStudy (anonymous):

or solve for x and plug into cos(x) cos(9x)=1/2 9x={cos^-1(1/2)+2*pi*C,-cos^-1(1/2)+2*pi*C} 9x={pi/3+2*pi*C,-pi/3+2*pi*C} cos(x)={cos(pi/27+2/9*pi*C),cos(-pi/27+2/9*pi*C)}

OpenStudy (venomblast):

confusing